# AAVENUE: Detecting LLM Biases on NLU Tasks in AAVE via a Novel Benchmark

### v1

@@ Abstract
Detecting biases in natural language understanding (NLU) for African American Vernacular English (AAVE) is crucial to developing inclusive natural language processing (NLP) systems. To address dialect-induced performance discrepancies, we introduce AAVENUE (AAVE Natural Language Understanding Evaluation), a benchmark for evaluating large language model (LLM) performance on NLU tasks in AAVE and Standard American English (SAE). AAVENUE builds upon and extends existing benchmarks like VALUE, replacing deterministic syntactic and morphological transformations with a more flexible methodology leveraging LLM-based translation with few-shot prompting, improving performance across our evaluation metrics when translating key tasks from the GLUE and SuperGLUE benchmarks. We compare AAVENUE and VALUE translations using five popular LLMs and a comprehensive set of metrics including fluency, BARTScore, quality, coherence, and understandability. Additionally, we recruit fluent AAVE speakers to validate our translations for authenticity. Our evaluations reveal that LLMs consistently perform better on SAE tasks than AAVE-translated versions, underscoring inherent biases and highlighting the need for more inclusive NLP models. We have open-sourced our source code on GitHub and created a website to showcase our work at https://aavenue.live.
@@
Abhay Guptathanks: Lead Author           Philip Meng           Ece Yurtseven           Sean O’Brien thanks: Senior Author          Kevin Zhu22footnotemark: 2 Algoverse AI Research abhaygupta1266@gmail.com, kevin@algoverse.us
@@ 2 Dataset
To evaluate the performance of models on SAE and AAVE, we selected five key tasks from the GLUE and SuperGLUE benchmarks, knwon for their diversity and relevance to natural language understanding tasks (Wang et al., 2019, 2020). These tasks include BoolQ, MultiRC, SST-2, COPA, and WSC.
@@ 2.3.1 Translation Metrics Analysis
The evaluation of our AAVE translations against those from the VALUE benchmark demonstrates clear advantages across several key metrics. (1) Quality: Our translations scored higher in quality, with our BoolQ passages receiving a score of 76.57 compared to VALUE’s 58.21. This shows that our translations are better in terms of overall quality, including accuracy, style, and appropriateness. (see Table 8). (2) Fluency: Our translations achieved a fluency score of 70.72 in BoolQ passages, compared to VALUE’s 57.32. This suggests that our translations are better written and more grammatically sound, resulting in improved readability. (see Table 6). (3) Coherence: Our translations exhibited better coherence, with scores like 74.39 for BoolQ passages versus VALUE’s 62.10, showing that our translations make more logical sense and maintain consistency throughout the text (see Table 7). (4) Understandability: In terms of understandability, our translations were clearer and more accessible, scoring 76.53 in BoolQ passages compared to 52.26 for VALUE. This indicates that our translations are easier for readers to understand (see Table 9). (5) BARTScores: Finally, our translations achieved better BARTScores, with a score of -1.44 in BoolQ passages compared to VALUE’s -1.54, indicating that our translations are closer to human-produced texts and better maintain fidelity to the original content (see Table 10). Collectively, these metrics confirm that our translations consistently outperform those from the VALUE benchmark, providing superior quality, fluency, coherence, understandability, and fidelity.
@@ 2.3.2 Comparison Scores Analysis
The comparison scores provide a direct evaluation of our AAVE translations against those from the VALUE benchmark across various tasks and models. As shown in Tables 4 and 5, our AAVE translations were consistently preferred over the VALUE translations by the LLMs we evaluated. For instance, in the BoolQ passage task using GPT-4-turbo, our translations were preferred 94.51% of the time compared to VALUE’s 4.62% (see Table 4). Similarly, in the COPA premise task, GPT-4o-mini showed a preference for our translations 90.42% of the time, with VALUE translations receiving only 9.38% (see Table 4). When evaluated with the Gemini-1.5-flash model on the BoolQ passage task, our translations were preferred 93.85% of the time compared to VALUE’s 6.15% (see Table 5). These consistent preferences across multiple models and tasks show the effectiveness of our translations in capturing the nuances of AAVE while maintaining the meaning of the original text. The results indicate that our approach to translating SAE to AAVE not only meets but often exceeds the standards set by the VALUE benchmark.
@@ 2.4 Validation by AAVE Speakers
We recruited 10 fluent AAVE speakers from the Bronx and Queens area to assess the quality of our AAVE translations across five tasks on a scale of 1 to 10, specifically rating how accurately the translations reflect AAVE. The results, detailed in Tables 14 and 15, show that the average scores across the tasks ranged from 7.02 to 7.27. The highest average score was observed for MultiRC (7.27), while BoolQ received the lowest average score (7.02). These consistent ratings across multiple fluent speakers, reinforce the reliability of our AAVE translations. The slightly higher scores for subsets like WSC and MultiRC suggest that our translations were particularly effective in maintaining clarity and coherence in more complex linguistic structures. Overall, the validators’ feedback confirms the quality and authenticity of our translations, aligning well with our evaluation metrics and further validating our approach.
@@ 3 Results
3.1 Accuracy Scores Across Tasks
We evaluated the accuracy of the translations in five tasks using five LLMs. The accuracy scores show the performance of each model in SAE and AAVE translations, as shown in Tables 1 and 2.
Table 1: Accuracy Scores for GPT Models (%)
All scores are presented in the format SAE/AAVE.
Task GPT-4o-mini GPT-4-turbo GPT-4o
SST-2 90.40/88.40 (-2.0) 94.00/92.80 (-1.2) 88.80/87.30 (-1.5)
BoolQ 88.29/85.29 (-3.0) 88.09/86.49 (-1.6) 89.19/86.89 (-2.3)
COPA 95.40/93.20 (-2.2) 97.60/96.80 (-0.8) 97.20/96.40 (-0.8)
WSC 60.03/57.90 (-2.1) 69.60/68.69 (-0.9) 70.36/67.02 (-3.3)
MultiRC 84.50/72.00 (-12.5) 86.20/73.70 (-12.5) 87.50/71.30 (-16.2)
Table 2: Accuracy Scores for Gemini Models (%)
All scores are presented in the format SAE/AAVE.
Task Gemini-1.5-Flash Gemini-1.5-Pro
SST-2 87.70/87.10 (-0.6) 92.00/91.40 (-0.6)
BoolQ 89.69/87.29 (-2.4) 89.49/85.89 (-3.6)
COPA 91.40/92.00 (+0.6) 97.40/95.80 (-1.6)
WSC 48.78/48.48 (-0.3) 51.37/51.22 (-0.2)
MultiRC 84.10/70.70 (-13.4) 85.90/71.90 (-14.0)
@@ 3.2 Accuracy Score Analysis
The accuracy scores from Tables 1 and 2 highlight consistent performance drops when handling AAVE translations across all models. MultiRC and WSC exhibited the largest declines, indicating challenges in complex reading comprehension and pronoun resolution tasks. GPT-4-turbo generally showed smaller accuracy drops, suggesting better adaptation to AAVE, while other models like GPT-4o-mini struggled more, particularly with contextually demanding tasks. Overall, these results underscore the need for more inclusive training data and models better equipped to handle AAVE.
@@ 3.3 Intersection Over Union (IoU) Analysis
The IoU table shows the percentage of incorrectly answered questions in both our AAVE translation and SAE across five models and tasks. The results, shown in Tables 12 and 13 in the Appendix, indicate minimal overlap in incorrect responses, suggesting challenges in handling each dialect. The analysis reveals that challenges are dialect-specific, as there was minimal overlap in errors between SAE and AAVE. However, the high IoU in WSC for models like Gemini-1.5-Flash indicates that some tasks present had significant difficulties in both dialects. These results show the importance of developing more robust models that can handle the distinct features of AAVE, as current models show variability in managing dialect-specific errors.
@@ 6 Limitations
We are mindful of the ethical implications of our research, which focuses on evaluating and addressing dialectal biases in LLMs through the development of the AAVENUE benchmark. While some data used in this study is publicly available, we also collected original data with careful consideration to ensure cultural and linguistic authenticity. The data collection process adhered to ethical guidelines, and all participants provided informed consent. Fluent AAVE speakers were recruited for the validation process to ensure the translations accurately reflect cultural and linguistic nuances. Throughout our research, we aimed to avoid potential harm and bias, with the goal of contributing to the development of more inclusive NLP systems. We have made efforts to report our findings transparently and objectively. We believe our research advances the field while adhering to rigorous ethical standards.
Appendix A Appendix
A.1 Few-Shot AAVE Translation from VALUE
Table 3: Sample Few-Shot Prompted AAVE Translations from VALUE
AAVE Translation
@@ A.1 Few-Shot AAVE Translation from VALUE
A.2 Detailed Comparison Scores
Table 4: LLM-Based Binary Preference Scores Across Tasks - GPT Models (AAVE/VALUE/About the Same)
Task GPT-4o-mini GPT-4-turbo GPT-4o
BoolQ (P) 88.79/10.33/0.88 94.51/4.62/0.88 94.95/4.84/0.22
BoolQ (Q) 82.33/17.14/0.53 78.13/20.08/1.58 80.34/18.72/0.84
MultiRC (P) 85.71/14.29/0.00 71.43/28.57/0.00 100.00/0.00/0.00
MultiRC (Q) 66.09/28.16/5.75 62.07/32.18/5.75 74.14/20.69/4.60
SST-2 87.25/9.92/2.83 80.88/12.04/7.08 87.39/9.92/2.69
COPA (P) 90.42/9.38/0.21 72.92/26.88/0.21 85.42/14.17/0.42
COPA (C1) 78.46/20.68/0.85 67.38/31.13/1.49 76.33/22.39/1.28
COPA (C2) 77.23/22.34/0.43 64.26/33.62/2.13 79.15/20.21/0.64
WSC (P) 83.80/16.20/0.00 73.99/26.01/0.00 86.76/13.08/0.16
Table 5: LLM-Based Binary Preference Scores Across Tasks - Gemini Models (AAVE/VALUE/About the Same)
Task Gemini-1.5-flash Gemini-1.5-pro
BoolQ (P) 93.85/6.15/0.00 88.13/10.99/0.88
BoolQ (Q) 75.18/24.40/0.42 44.16/53.42/2.42
MultiRC (P) 42.86/57.14/0.00 14.29/85.71/0.00
MultiRC (Q) 56.90/39.66/3.45 44.25/46.55/9.20
SST-2 84.70/12.89/2.41 72.24/20.40/7.37
COPA (P) 75.21/24.58/0.21 65.83/33.75/0.42
COPA (C1) 69.51/29.42/1.07 58.00/39.23/2.77
COPA (C2) 68.51/31.28/0.21 56.60/40.64/2.77
WSC (P) 84.11/15.26/0.62 57.17/42.21/0.62
A.3 Fluency Scores
Table 6: Fluency Scores Across Tasks
Task AAVE (GPT-4o-mini) VALUE
BoolQ (Passages) 70.7169 57.3199
BoolQ (Questions) 54.0808 52.8357
COPA (Choice 1) 58.8326 51.5955
COPA (Choice 2) 55.5880 54.7378
COPA (Premise) 54.9623 60.1386
MultiRC (Questions) 66.6923 59.2436
MultiRC (Paragraphs) 51.7319 57.6290
SST-2 (Sentences) 71.6592 64.7925
WSC (Paragraphs) 64.9487 63.2925
A.4 Coherence Scores
Table 7: Coherence Scores Across Tasks
Task AAVE (GPT-4o-mini) VALUE Translations
BoolQ (P) 74.3910 62.0983
BoolQ (Q) 56.2019 51.9285
MultiRC (P) 74.6919 65.1029
MultiRC (Q) 60.1021 56.1028
SST-2 83.1921 74.1029
COPA (P) 71.2058 63.5912
COPA (C1) 65.0852 54.5192
COPA (C2) 56.2019 51.2912
WSC (P) 64.0991 61.2017
A.5 Quality Scores
Table 8: Quality Scores Across Tasks
Task AAVE (GPT-4o-mini) VALUE Translations
BoolQ (P) 76.5699 58.2102
BoolQ (Q) 64.9691 56.7069
MultiRC (P) 64.9026 53.2962
MultiRC (Q) 62.7819 55.2493
SST-2 70.1046 66.9014
COPA (P) 67.6039 66.4782
COPA (C1) 73.5693 67.2891
COPA (C2) 75.2962 61.8783
WSC (P) 66.1021 66.3019
A.6 Understandability Scores
Table 9: Understandability Scores Across Tasks
Task AAVE (GPT-4o-mini) VALUE Translations
BoolQ (P) 76.5327 52.2594
BoolQ (Q) 64.4789 51.5689
MultiRC (P) 66.7032 57.1391
MultiRC (Q) 68.2280 53.2496
SST-2 82.5961 47.6228
COPA (P) 64.3259 57.6248
COPA (C1) 59.3256 48.1259
COPA (C2) 63.5832 57.3462
WSC (P) 76.7991 53.6917
A.7 BARTScores
Table 10: BARTScores Across Tasks
Task AAVE (GPT-4o-mini) VALUE Translations
BoolQ (P) -1.4402 -1.5352
BoolQ (Q) -1.6750 -2.8862
MultiRC (P) -1.8827 -1.7628
MultiRC (Q) -1.8490 -2.4334
SST-2 -2.5356 -2.5991
COPA (P) -1.8744 -3.0585
COPA (C1) -1.9161 -3.1870
COPA (C2) -2.0007 -3.1323
WSC (P) -1.9589 -2.5989
A.8 BLEU Scores
Table 11: BLEU Scores Across Tasks (Percentage Below Threshold)
Task < 0.7 < 0.5 < 0.3

### v2

@@ Abstract
Detecting biases in natural language understanding (NLU) for African American Vernacular English (AAVE) is crucial to developing inclusive natural language processing (NLP) systems. To address dialect-induced performance discrepancies, we introduce AAVENUE (AAVE Natural Language Understanding Evaluation), a benchmark for evaluating large language model (LLM) performance on NLU tasks in AAVE and Standard American English (SAE). AAVENUE builds upon and extends existing benchmarks like VALUE, replacing deterministic syntactic and morphological transformations with a more flexible methodology leveraging LLM-based translation with few-shot prompting, improving performance across several evaluation metrics when translating key tasks from the GLUE and SuperGLUE benchmarks. We compare AAVENUE and VALUE translations using five popular LLMs and a comprehensive set of metrics including fluency, BARTScore, quality, coherence, and understandability. Additionally, the fluency of AAVENUE is validated by annotations from AAVE speakers. Our evaluations reveal that LLMs consistently perform better on SAE tasks than AAVE-translated versions, underscoring inherent biases and highlighting the need for more inclusive NLP models. This work has been accepted at the NLP4I workshop at EMNLP 2024. We have open-sourced our source code on GitHub and created a website to showcase our work at https://aavenue.live.
@@
Abhay Guptathanks: Lead Author           Ece Yurtseven           Philip Meng           Sean O’Brienthanks: Senior Authors           Kevin Zhu22footnotemark: 2 Algoverse AI Research abhaygupta1266@gmail.com, kevin@algoverse.us
@@ 2 Dataset
To evaluate the performance of models on SAE and AAVE, we selected five key tasks from the GLUE and SuperGLUE benchmarks, known for their diversity and relevance to natural language understanding tasks (Wang et al., 2019, 2020). These tasks include BoolQ, MultiRC, SST-2, COPA, and WSC.
@@ 2.2 Translating SAE to AAVE
Task Quality Fluency Coherence Understandability BARTScore
AAVE VALUE AAVE VALUE AAVE VALUE AAVE VALUE AAVE VALUE
BoolQ (P) 76.57 58.21 70.72 57.32 74.39 62.10 76.53 52.26 -1.44 -1.54
BoolQ (Q) 64.97 56.71 54.08 52.84 56.20 51.93 64.48 51.57 -1.68 -2.89
MultiRC (P) 64.90 53.30 51.73 57.63 74.69 65.10 66.70 57.14 -1.88 -1.76
MultiRC (Q) 62.78 55.25 66.69 59.24 60.10 56.10 68.23 53.25 -1.85 -2.43
SST-2 70.10 66.90 71.66 64.79 83.19 74.10 82.60 47.62 -2.54 -2.60
COPA (P) 67.60 66.48 54.96 60.14 71.21 63.59 64.33 57.62 -1.87 -3.06
COPA (C1) 73.57 67.29 58.83 51.60 65.09 54.52 59.33 48.13 -1.92 -3.19
COPA (C2) 75.30 61.88 55.59 54.74 56.20 51.29 63.58 57.35 -2.00 -3.13
WSC (P) 66.10 66.30 64.95 63.29 64.10 61.20 76.80 53.69 -1.96 -2.60
Table 1: Comparison of Translation Metrics Across Tasks
Task GPT-4o-mini GPT-4-turbo GPT-4o Gemini-1.5-Flash Gemini-1.5-Pro
BoolQ (P) 88.79/10.33/0.88 94.51/4.62/0.88 94.95/4.84/0.22 93.85/6.15/0.00 88.13/10.99/0.88
BoolQ (Q) 82.33/17.14/0.53 78.13/20.08/1.58 80.34/18.72/0.84 75.18/24.40/0.42 44.16/53.42/2.42
MultiRC (P) 85.71/14.29/0.00 71.43/28.57/0.00 100.00/0.00/0.00 42.86/57.14/0.00 14.29/85.71/0.00
MultiRC (Q) 66.09/28.16/5.75 62.07/32.18/5.75 74.14/20.69/4.60 56.90/39.66/3.45 44.25/46.55/9.20
SST-2 87.25/9.92/2.83 80.88/12.04/7.08 87.39/9.92/2.69 84.70/12.89/2.41 72.24/20.40/7.37
COPA (P) 90.42/9.38/0.21 72.92/26.88/0.21 85.42/14.17/0.42 75.21/24.58/0.21 65.83/33.75/0.42
COPA (C1) 78.46/20.68/0.85 67.38/31.13/1.49 76.33/22.39/1.28 69.51/29.42/1.07 58.00/39.23/2.77
COPA (C2) 77.23/22.34/0.43 64.26/33.62/2.13 79.15/20.21/0.64 68.51/31.28/0.21 56.60/40.64/2.77
WSC (P) 83.80/16.20/0.00 73.99/26.01/0.00 86.76/13.08/0.16 84.11/15.26/0.62 57.17/42.21/0.62
Table 2: LLM-Based Binary Preference Scores Across Tasks for GPT and Gemini Models (AAVE/VALUE/About the Same)
@@ 2.3.1 Translation Metrics Analysis
The evaluation of our AAVE translations against those from the VALUE benchmark demonstrates clear advantages across several key metrics.
(1) Quality: Our translations scored higher in quality, with our BoolQ passages receiving a score of 76.57 compared to VALUE’s 58.21. This shows that our translations are better in terms of overall quality, including accuracy, style, and appropriateness.
(2) Fluency: Our translations achieved a fluency score of 70.72 in BoolQ passages, compared to VALUE’s 57.32. This suggests that our translations are better written and more grammatically sound, resulting in improved readability.
(3) Coherence: Our translations exhibited better coherence, with scores like 74.39 for BoolQ passages versus VALUE’s 62.10, showing that our translations make more logical sense and maintain consistency throughout the text.
(4) Understandability: In terms of understandability, our translations were clearer and more accessible, scoring 76.53 in BoolQ passages compared to 52.26 for VALUE. This indicates that our translations are easier for readers to understand.
(5) BARTScores: Finally, our translations achieved better BARTScores, with a score of -1.44 in BoolQ passages compared to VALUE’s -1.54, indicating that our translations are closer to human-produced texts and better maintain fidelity to the original content.
Collectively, these metrics confirm that our translations consistently outperform those from the VALUE benchmark, providing superior quality, fluency, coherence, understandability, and fidelity.
Task GPT-4o-mini GPT-4-turbo GPT-4o Gemini-1.5-Flash Gemini-1.5-Pro
SST-2 90.40/88.40 (-2.0) 94.00/92.80 (-1.2) 88.80/87.30 (-1.5) 87.70/87.10 (-0.6) 92.00/91.40 (-0.6)
BoolQ 88.29/85.29 (-3.0) 88.09/86.49 (-1.6) 89.19/86.89 (-2.3) 89.69/87.29 (-2.4) 89.49/85.89 (-3.6)
COPA 95.40/93.20 (-2.2) 97.60/96.80 (-0.8) 97.20/96.40 (-0.8) 91.40/92.00 (+0.6) 97.40/95.80 (-1.6)
WSC 60.03/57.90 (-2.1) 69.60/68.69 (-0.9) 70.36/67.02 (-3.3) 48.78/48.48 (-0.3) 51.37/51.22 (-0.2)
MultiRC 84.50/72.00 (-12.5) 86.20/73.70 (-12.5) 87.50/71.30 (-16.2) 84.10/70.70 (-13.4) 85.90/71.90 (-14.0)
Table 3: Accuracy Scores for GPT and Gemini Models (%) All scores are presented in the format SAE/AAVE.
Task GPT-4o-mini GPT-4-turbo GPT-4o Gemini-1.5-Flash Gemini-1.5-Pro
SST-2 8.40 5.10 9.80 10.40 6.20
BoolQ 10.21 10.71 8.91 8.51 8.41
COPA 3.00 1.60 2.00 5.80 1.80
WSC 35.56 24.01 25.68 49.54 44.53
MultiRC 9.60 9.00 8.30 9.90 7.90
Table 4: Intersection Over Union Between Incorrect Answers for SAE and AAVE Across Tasks (%)
@@ 2.3.2 Comparison Scores Analysis
The comparison scores provide a direct evaluation of our AAVE translations against those from the VALUE benchmark across various tasks and models. As shown in Table 2, our AAVE translations were consistently preferred over the VALUE translations by the LLMs we evaluated. For instance, in the BoolQ passage task using GPT-4-turbo, our translations were preferred 94.51% of the time compared to VALUE’s 4.62%. Similarly, in the COPA premise task, GPT-4o-mini showed a preference for our translations 90.42% of the time, with VALUE translations receiving only 9.38%. When evaluated with the Gemini-1.5-Flash model on the BoolQ passage task, our translations were preferred 93.85% of the time compared to VALUE’s 6.15%. These consistent preferences across multiple models and tasks demonstrate the effectiveness of our translations in capturing the nuances of AAVE while maintaining the meaning of the original text. The results indicate that our approach to translating SAE to AAVE not only meets but often exceeds the standards set by the VALUE benchmark.
@@ 2.4 Validation by AAVE Speakers
We recruited 10 fluent AAVE speakers from the Bronx and Queens area to assess the quality of our AAVE translations across five tasks on a scale of 1 to 10, specifically rating how accurately the translations reflect AAVE. The results, detailed in Table 6 in Section B, show that the average scores across the tasks ranged from 7.02 to 7.27. The highest average score was observed for MultiRC (7.27), while BoolQ received the lowest average score (7.02). These consistent ratings across multiple fluent speakers reinforce the reliability of our AAVE translations. The slightly higher scores for tasks like WSC and MultiRC suggest that our translations were particularly effective in maintaining clarity and coherence in more complex linguistic structures. Overall, the validators’ feedback confirms the quality and authenticity of our translations, aligning well with our evaluation metrics and further validating our approach.
@@ 3 Results
We evaluated the accuracy of the translations in five tasks using five LLMs. The accuracy scores show the performance of each model in SAE and AAVE translations, highlighting consistent performance drops when handling AAVE translations across all models, as shown in Table 3.
@@ 3.1 Accuracy Score Analysis
The accuracy scores from Table 3 highlight consistent performance drops when handling AAVE translations across all models. MultiRC and WSC exhibited the largest declines, indicating challenges in complex reading comprehension and pronoun resolution tasks. GPT-4-turbo generally showed smaller accuracy drops, suggesting better adaptation to AAVE, while other models like GPT-4o-mini struggled more, particularly with contextually demanding tasks. Overall, these results underscore the need for more inclusive training data and models better equipped to handle AAVE.
@@ 3.2 Intersection Over Union (IoU) Analysis
The IoU table shows the percentage of incorrectly answered questions in both our AAVE translation and SAE across five models and tasks. The results, shown in Table 4, indicate minimal overlap in incorrect responses, suggesting challenges in handling each dialect. The analysis reveals that challenges are dialect-specific, as there was minimal overlap in errors between SAE and AAVE. However, the high IoU in WSC for models like Gemini-1.5-Flash indicates that some tasks present had significant difficulties in both dialects. These results show the importance of developing more robust models that can handle the distinct features of AAVE, as current models show variability in managing dialect-specific errors.
@@ 6 Limitations
We are mindful of the ethical implications of our research, which focuses on evaluating and addressing dialectal biases in LLMs through the development of the AAVENUE benchmark. While some data used in this study is publicly available, we also collected original data with careful consideration to ensure cultural and linguistic authenticity. The data collection process adhered to ethical guidelines, and all participants provided informed consent. Our human validators, who are fluent AAVE speakers from the Bronx, NY, were recruited to ensure the translations accurately reflect cultural and linguistic nuances. Annotators were compensated for their time and effort, and we encouraged them to take breaks if they felt overwhelmed during the annotation process. Throughout our research, we aimed to avoid potential harm and bias, with the goal of contributing to the development of more inclusive NLP systems. We have made efforts to report our findings transparently and objectively. We believe our research advances the field while adhering to rigorous ethical standards.
Appendix A Few-Shot AAVE Translation Examples
AAVE Translation Examples
@@ Appendix A Few-Shot AAVE Translation Examples
Table 5: Examples of the Few-Shot Prompted AAVE Translations from VALUE used in our experiments.
Appendix B Human Validators’ Scores
Validator COPA BoolQ MultiRC SST-2 WSC
Validator 1 6.9 7.1 7.4 6.6 7.3
Validator 2 7.0 7.4 7.5 7.4 8.3
Validator 3 7.6 6.5 7.0 7.2 6.9
Validator 4 6.5 6.9 7.0 6.6 7.0
Validator 5 7.5 6.9 7.4 7.0 6.9
Validator 6 7.3 7.0 7.2 7.3 7.2
Validator 7 7.6 7.2 7.5 7.5 7.1
Validator 8 7.2 7.1 7.3 7.1 7.4
Validator 9 7.3 7.2 7.4 7.3 7.3
Validator 10 7.3 6.9 7.0 6.9 7.1
Average 7.22 7.02 7.27 7.09 7.25
Table 6: Human Validators’ Scores for AAVE Translations Across All Tasks (Out of 10)
Appendix C BLEU Scores
Task BLEU <<< 0.7 (%) BLEU <<< 0.5 (%) BLEU <<< 0.3 (%)